OpenStudy (anonymous):

One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

OpenStudy (anonymous):

humans 50 and horses 24

OpenStudy (anonymous):

let no of horses = x and no of humans = y both horses and humans have one head each => x + y = 74 => x = 74 - y ------ (i) humans have two legs each and horses have four legs each => 4x + 2y = 196 substitute value of x from (i) => 4(74 - y) + 2y = 196 => 296 - 4y + 2y = 196 => 296 - 196 = 2y => 100 = 2y => 50 = y so x = 74 - 50 = 24 so there are 24 horses and 50 humans